# Camera: Multi-Matrix Joint Compression for MoE Models via Micro-Expert Redundancy Analysis

### v3

@@ Abstract
Large Language Models (LLMs) with Mixture-of-Experts (MoE) architectures are distinguished by their strong performance scaling with increasing parameters across a wide range of tasks, yet they also suffer from substantial computational and storage overheads. Notably, the performance gains of MoE models do not scale proportionally with the growth in expert parameters. While prior works attempt to reduce parameters via expert-level pruning, merging, or decomposition, they still suffer from challenges in both performance and computational efficiency. In this paper, we address these challenges by introducing micro-expert as a finer-grained compression unit that spans across matrices. We first establish a more fundamental perspective, viewing MoE layers as mixtures of micro-experts, and present Camera, a lightweight and training-free framework for identifying micro-expert redundancy. Our analysis uncovers significant variance in micro-expert contributions during decoding. Based on this insight, we further propose Camera-P, a structured micro-expert pruning framework, and Camera-Q, a mixed-precision quantization idea designed for micro-experts. Extensive experiments on nine downstream tasks show that Camera-P consistently outperforms strong baselines under pruning ratios ranging from 20% to 60%. Furthermore, Camera-Q achieves superior results under aggressive 2-bit quantization, surpassing existing matrix- and channel-level ideas. Notably, our method enables complete micro-expert analysis of Qwen2-57B-A14B in less than 5 minutes on a single NVIDIA A100-40GB GPU.
@@ 1 Introduction
Large Language Models (LLMs) based on Mixture-of-Experts (MoE) architecture leverage sparse Feed-Forward Network (FFN) structures to enable efficient model scaling, where each time only a subset of experts is activated by a router (Shazeer et al. 2017; Fedus, Zoph, and Shazeer 2022). This design facilitates the scaling of LLMs to hundreds of billions of parameters. Many well-known open-source models, such as Qwen3-MoE (Yang et al. 2025), Deepseek-MoE (Liu et al. 2024a), Kimi-K2 (Team et al. 2025), and Mixtral-MoE (Jiang et al. 2024), adopt the MoE architecture and achieve impressive results on downstream tasks. However, such substantial expansion in parameters does not bring proportional improvements in model capability, while also coupling with prohibitive computational and storage overheads—highlighting the structural redundancy inherent in MoE designs.
Pruning is a widely used strategy to reduce redundancy in MoE models. Existing methods for reducing expert parameters primarily fall into two categories: expert pruning and expert merging. Expert pruning removes either all or part of the parameters within each expert (Lu et al. 2024; Yang et al. 2024). Full pruning inevitably results in severe information loss, while partial pruning often yields suboptimal performance due to the lack of reliable measures for intra-expert importance. Expert merging, on the other hand, seeks to mitigate information loss by assuming functional similarity among experts (Yang et al. 2024; Li et al. 2024). However, this assumption rarely holds in practice, thereby limiting the effectiveness of merging-based strategies. An emerging direction exemplified by $D^{2}$-MoE attempts to relax this assumption by combining expert merging with delta compression. Specifically, it constructs a shared base weight and expert-specific delta weights, both compressed using low-rank approximation (Gu et al. 2025; Ai et al. 2025).
@@
Although recent studies suggest that partial expert pruning is more effective than full expert removal (Xie et al. 2024), especially in modern MoE models with a large number of experts, challenges remain due to limited understanding of expert behavior within MoE layers. In particular, it is difficult to identify essential parameters and determine optimal retention ratios for each expert. A further limitation of most existing methods is that they compress each matrix in the expert independently, overlooking the functional dependencies among different matrices. To shed light on the internal mechanism of MoE layers, we introduce a finer-grained structural perspective by viewing each MoE layer as a mixture of micro-experts, where each micro-expert is jointly defined by the three transformations—up_proj, gate_proj and down_proj, as shown in Figure 1. Our analysis reveals that the output of an MoE layer is a linear combination of all micro-experts, whose relative importance varies significantly. This insight forms the critical basis for the compression strategies proposed in this work.
@@ 2.1 Model Compression
This work focuses on model weight compression. The primary approaches for compressing LLM weights include quantization, pruning, knowledge distillation, and low-rank approximation. Most quantization methods target post-training conversion of well-trained model weights into low-bit representations (Frantar et al. 2022; Xiao et al. 2023). The state-of-the-art quantization algorithms can reduce weights to 3-bit precision with negligible performance degradation (Liu et al. 2025; Shao et al. 2024). With the aid of carefully designed mixed-precision strategies, even sub-3-bit quantization can maintain performance comparable to the original model (Tseng et al. 2024; Xu et al. 2025). Pruning, on the other hand, reduces model size by eliminating less critical weights. Unstructured pruning offers maximum flexibility in weight removal and typically results in the least performance drop (Sun et al. 2024). Structured pruning imposes constraints on which weights can be pruned based on their location or structure, which may lead to slightly larger performance loss but enable actual inference speedup and make it more suitable for deployment (Frantar and Alistarh 2023). Knowledge distillation guided by loss minimization (Xu et al. 2024), and low-rank approximation typically via singular value decomposition (Ping et al. 2024), are also commonly integrated into broader compression frameworks. For further reading, please refer to survey (Zhou et al. 2024). As MoE becomes a dominant architecture for scaling LLMs, several compression techniques are developed specifically for MoE-based models.
@@ 2.2 Pruning Methods for MoE Models
Most pruning methods for MoE models focus on expert-level compression rather than on individual matrices. These fall mainly into two categories: direct pruning and expert merging (sometimes followed by pruning). Direct pruning methods like NAEE exhaustively evaluate expert combinations and retain the one yielding the lowest loss on calibration data (Lu et al. 2024). While effective on Mixtral-MoE, this brute-force approach does not scale well to modern sparse MoEs. MoE-$I^{2}$ (Yang et al. 2024) and MoE-Pruner (Xie et al. 2024) partially prune expert weights, but struggle to balance identifying important weights and achieving speedup. Expert merging methods fall into two subgroups. The first assumes functional similarity across experts, grouping and merging them based on activation frequency, output similarity, or knowledge distribution (e.g., MC-SMoE (Li et al. 2024), HC-SMoE (Chen et al. 2024), TAP (Zhang et al. 2024), EEP (Liu et al. 2024b)). This idealized assumption often limits performance. The second extracts common components from all experts (e.g., $D^{2}$-MoE (Gu et al. 2025), ResMoE (Ai et al. 2025), Sub-MoE (Li et al. 2025a), MoE-SVD (Li et al. 2025b)), then applies low-rank compression to approximate residuals. However, these approaches overlook the functional integrity across the three transformations in FFNs and often fail to preserve the most critical parameters. Our method, Camera-P, belongs to the partial pruning family but differs in two key ways. First, it performs fine-grained pruning at the micro-expert level, capturing cross-matrix coordination in FFNs. Second, it is powered by Camera, our fast and accurate importance estimator, addressing a major bottleneck of existing methods.
@@ 2.3 Quantization Methods for MoE Models
Most quantization work targets LLMs rather than MoEs, with MoE studies focusing mainly on assigning different bit-widths across experts. Studies like MC (Huang et al. 2025), MxMoE (Duanmu et al. 2025), and AFGQ (Xie et al. 2025) assign bit-widths based on a combination of activation frequency and weight sensitivity. In contrast, our Camera-Q is built on Camera, requiring no activation statistics, no pre-quantization, and no expensive evolutionary search.
@@ 4.1 Pruning Framework Camera-P
4.2 Experimental Setup
Models and Data
Baselines
We compare Camera-P with two strong baselines: NAEE (Lu et al. 2024), which prunes entire experts directly, and $D^{2}$-MoE (Gu et al. 2025), a recent strong merge-then-compress approach. To address the scalability issue with $N_{E}\geq 8$ in NAEE, we follow MoE-$I^{2}$ (Yang et al. 2024) and adopt efficient genetic search. Camera-P uses $\alpha=0.95,0.95,1.00$ for the three models.
Evaluation
To evaluate baseline performance, we calculate perplexity on randomly sampled sequences from Wikitext2 (Merity et al. 2017) and C4 (Raffel et al. 2020)—lower is better. We also report accuracy on nine zero-shot downstream tasks, including Winogrande (Sakaguchi et al. 2021), HellaSwag (Zellers et al. 2019), PIQA (Bisk et al. 2020), BoolQ (Clark et al. 2019), ARC-e/ARC-c (Clark et al. 2018), OBQA (Mihaylov et al. 2018), MathQA (Amini et al. 2019) and RTE (Wang et al. 2018). We prioritize outputting acc_norm from LM-Evaluation-Harness [^1] https://github.com/EleutherAI/lm-evaluation-harness.
4.3 Main Results
@@
Qwen3-30B-A3B (128 common experts, top-8 experts are activated)
@@ A.1 Derivation of Micro-Expert
$\mathbf{y}$ $=\sum_{i}^{N_{E}}A_{i}\cdot\sum_{j}^{d_{\mathrm{ff}}}\sigma\left(\mathbf{W}_{i}^{\mathrm{up}}[j,:]\mathbf{x}\right)\cdot\mathbf{W}_{i}^{\mathrm{gate}}[j,:]\mathbf{x}\cdot\mathbf{W}_{i}^{\mathrm{down}}[:,j].$
@@
$\mathbf{y}$ $=\sum_{i}^{N_{E}\times d_{\mathrm{ff}}}A_{i}\cdot\sigma\left(\mathbf{w}_{i}^{\mathrm{up}}\mathbf{x}\right)\cdot\mathbf{w}_{i}^{\mathrm{gate}}\mathbf{x}\cdot\mathbf{w}^{\mathrm{down}}.$
We separate the scalar terms in the above equation and denote them as $\phi_{i}=A_{i}\cdot\sigma\left(\mathbf{w}_{i}^{\mathrm{up}}\mathbf{x}\right)\cdot\mathbf{w}_{i}^{\mathrm{gate}}\mathbf{x}$. Hence, we have:
@@ A.3 Supplementary Micro-Expert Distributions
Figure 6: Distributions of micro-experts within each expert based on global ranking from Camera on Deepseek-MoE-16B. We list all 66 experts, where ‘S0/S1’ denotes the shared experts, and the rest are non-shared common experts. For 6(c), we present the results of the pruned model at $\lambda=40\%$, while the other subfigures correspond to $\lambda=20\%$.
@@
In Section 6.1 of the main text, we presented the micro-expert analysis of layer 12 in the Deepseek-MoE-16B model with a overall pruning ratio of $\lambda=20\%$. Here, we further supplement the analysis with results from layers 3 and 28, as shown in Figures 6(a) and 6(b). First, we observe that the importance of the first two shared experts, S0 and S1, is quite similar across layers, though not exactly the same. In contrast, the importance of the remaining experts varies significantly. Additionally, we also find that the shared experts tend to be less important in the lower layers (e.g., layers 2 and 3), as illustrated in Figure 6(a). Finally, we highlight a seemingly important property of the Camera algorithm: the micro-expert analysis results—namely, the importance rankings—remain largely consistent across different values of $\lambda$. As shown in Figure 6(c), when $\lambda=40\%$, the analysis results closely resemble those in Figure 2.
@@
As shown in Figure 7, the micro-expert distribution of the Qwen3-30B-A3B model is similar to that of Deepseek-MoE-16B, exhibiting large disparities in expert importance. Unlike Deepseek-MoE-16B, Qwen3-30B-A3B does not employ shared experts—all 128 experts are treated equally. Moreover, we observe that Qwen3-30B-A3B appears to have a higher proportion of unimportant experts. As shown in Figures 7(b) and 7(c), more expert box plots are concentrated near the top. We speculate that this may result from insufficient weight utilization, suggesting potential for continued training and further knowledge injection.

### v4

@@ Abstract
Large Language Models (LLMs) with Mixture-of-Experts (MoE) architectures are distinguished by their strong performance scaling with increasing parameters across a wide range of tasks, yet they also suffer from substantial computational and storage overheads. Notably, the performance gains of MoE models do not scale proportionally with the growth in expert parameters. While prior works attempt to reduce parameters via expert-level pruning, merging, or decomposition, they still suffer from challenges in both performance and computational efficiency. In this paper, we address these challenges by introducing micro-expert as a finer-grained compression unit that spans across matrices. We first establish a more fundamental perspective, viewing MoE layers as mixtures of micro-experts, and present Camera, a lightweight and training-free framework for identifying micro-expert redundancy. Our analysis uncovers significant variance in micro-expert contributions during decoding. Based on this insight, we further propose Camera-P, a structured micro-expert pruning framework, and Camera-Q, a mixed-precision quantization idea designed for micro-experts. Extensive experiments on nine downstream tasks show that Camera-P consistently outperforms strong baselines under pruning ratios ranging from 20% to 60%. Furthermore, Camera-Q achieves superior results under aggressive 2-bit quantization, surpassing existing matrix- and channel-level ideas. Notably, our method enables complete micro-expert analysis of Qwen2-57B-A14B in less than 5 minutes on a single NVIDIA A100-40GB GPU. Code is available at https://github.com/xuyuzhuang11/CAMERA.
@@ 1 Introduction
Large Language Models (LLMs) based on Mixture-of-Experts (MoE) architecture leverage sparse Feed-Forward Network (FFN) structures to enable efficient model scaling, where each time only a subset of experts is activated by a router (moe2017; switch2022). This design facilitates the scaling of LLMs to hundreds of billions of parameters. Many well-known open-source models, such as Qwen3-MoE (qwen2025), Deepseek-MoE (deepseek2024), Kimi-K2 (kimi2025), and Mixtral-MoE (mixtral2024), adopt the MoE architecture and achieve impressive results on downstream tasks. However, such substantial expansion in parameters does not bring proportional improvements in model capability, while also coupling with prohibitive computational and storage overheads—highlighting the structural redundancy inherent in MoE designs.
@@
Pruning is a widely used strategy to reduce redundancy in MoE models. Existing methods for reducing expert parameters primarily fall into two categories: expert pruning and expert merging. Expert pruning removes either all or part of the parameters within each expert (naee2024; moei2024). Full pruning inevitably results in severe information loss, while partial pruning often yields suboptimal performance due to the lack of reliable measures for intra-expert importance. Expert merging, on the other hand, seeks to mitigate information loss by assuming functional similarity among experts (moei2024; mcsmoe2023). However, this assumption rarely holds in practice, thereby limiting the effectiveness of merging-based strategies. An emerging direction exemplified by $D^{2}$-MoE attempts to relax this assumption by combining expert merging with delta compression. Specifically, it constructs a shared base weight and expert-specific delta weights, both compressed using low-rank approximation (delta2025; resmoe2025).
Although recent studies suggest that partial expert pruning is more effective than full expert removal (pruner2024), especially in modern MoE models with a large number of experts, challenges remain due to limited understanding of expert behavior within MoE layers. In particular, it is difficult to identify essential parameters and determine optimal retention ratios for each expert. A further limitation of most existing methods is that they compress each matrix in the expert independently, overlooking the functional dependencies among different matrices. To shed light on the internal mechanism of MoE layers, we introduce a finer-grained structural perspective by viewing each MoE layer as a mixture of micro-experts, where each micro-expert is jointly defined by the three transformations—up_proj, gate_proj and down_proj, as shown in Figure 1. Our analysis reveals that the output of an MoE layer is a linear combination of all micro-experts, whose relative importance varies significantly. This insight forms the critical basis for the compression strategies proposed in this work.
@@ 2.1 Model Compression
This work focuses on model weight compression. The primary approaches for compressing LLM weights include quantization, pruning, knowledge distillation, and low-rank approximation. Most quantization methods target post-training conversion of well-trained model weights into low-bit representations (gptq2022; smooth2023). The state-of-the-art quantization algorithms can reduce weights to 3-bit precision with negligible performance degradation (spinquant2025; omniquant2023). With the aid of carefully designed mixed-precision strategies, even sub-3-bit quantization can maintain performance comparable to the original model (quip2024; crvq2025). Pruning, on the other hand, reduces model size by eliminating less critical weights. Unstructured pruning offers maximum flexibility in weight removal and typically results in the least performance drop (wanda2024). Structured pruning imposes constraints on which weights can be pruned based on their location or structure, which may lead to slightly larger performance loss but enable actual inference speedup and make it more suitable for deployment (sparsegpt2023). Knowledge distillation guided by loss minimization (onebit2024), and low-rank approximation typically via singular value decomposition (delta2024), are also commonly integrated into broader compression frameworks. For further reading, please refer to survey (survey2024). As MoE becomes a dominant architecture for scaling LLMs, several compression techniques are developed specifically for MoE-based models.
@@ 2.2 Pruning Methods for MoE Models
Most pruning methods for MoE models focus on expert-level compression rather than on individual matrices. These fall mainly into two categories: direct pruning and expert merging (sometimes followed by pruning). Direct pruning methods like NAEE exhaustively evaluate expert combinations and retain the one yielding the lowest loss on calibration data (naee2024). While effective on Mixtral-MoE, this brute-force approach does not scale well to modern sparse MoEs. MoE-$I^{2}$ (moei2024) and MoE-Pruner (pruner2024) partially prune expert weights, but struggle to balance identifying important weights and achieving speedup. Expert merging methods fall into two subgroups. The first assumes functional similarity across experts, grouping and merging them based on activation frequency, output similarity, or knowledge distribution (e.g., MC-SMoE (mcsmoe2023), HC-SMoE (hcsmoe2024), TAP (tap2024), EEP (eep2024)). This idealized assumption often limits performance. The second extracts common components from all experts (e.g., $D^{2}$-MoE (delta2025), ResMoE (resmoe2025), Sub-MoE (submoe2025), MoE-SVD (moesvd2025)), then applies low-rank compression to approximate residuals. However, these approaches overlook the functional integrity across the three transformations in FFNs and often fail to preserve the most critical parameters. Our method, Camera-P, belongs to the partial pruning family but differs in two key ways. First, it performs fine-grained pruning at the micro-expert level, capturing cross-matrix coordination in FFNs. Second, it is powered by Camera, our fast and accurate importance estimator, addressing a major bottleneck of existing methods.
@@ 2.3 Quantization Methods for MoE Models
Most quantization work targets LLMs rather than MoEs, with MoE studies focusing mainly on assigning different bit-widths across experts. Studies like MC (mc2025), MxMoE (mxmoe2025), and AFGQ (afgq2025) assign bit-widths based on a combination of activation frequency and weight sensitivity. In contrast, our Camera-Q is built on Camera, requiring no activation statistics, no pre-quantization, and no expensive evolutionary search.
@@ 4.1 Pruning Framework Camera-P
Qwen3-30B-A3B (128 common experts, top-8 experts are activated)
@@
4.2 Experimental Setup
Models and Data
Baselines
We compare Camera-P with two strong baselines: NAEE (naee2024), which prunes entire experts directly, and $D^{2}$-MoE (delta2025), a recent strong merge-then-compress approach. To address the scalability issue with $N_{E}\geq 8$ in NAEE, we follow MoE-$I^{2}$ (moei2024) and adopt efficient genetic search. Camera-P uses $\alpha=0.95,0.95,1.00$ for the three models.
Evaluation
To evaluate baseline performance, we calculate perplexity on randomly sampled sequences from Wikitext2 (wiki22016) and C4 (C42020)—lower is better. We also report accuracy on nine zero-shot downstream tasks, including Winogrande (winogrande2021), HellaSwag (hellaswag2019), PIQA (piqa2020), BoolQ (boolq2019), ARC-e/ARC-c (arc2018), OBQA (obqa2018), MathQA (mathqa2019) and RTE (glue2018). We prioritize outputting acc_norm from LM-Evaluation-Harness [^1] https://github.com/EleutherAI/lm-evaluation-harness.
4.3 Main Results
@@ 6.3 Ablation on Balance Coefficient
Method Wiki2 C4 BoolQ OBQA RTE Wino. Hella. PIQA MathQA ARC-e ARC-c Avg.
openPangu Embedded-7B-V1.1 (dense model, only 1 expert)
Original (16-bit) 34.94 43.73 77.74 29.80 67.87 55.25 59.98 68.72 26.20 54.97 35.49 52.89
Camera-P (20%) 44.35 48.67 75.20 31.00 65.34 54.14 54.79 66.43 23.75 52.90 35.41 51.00
+Wanda (20%+20%) 42.96 48.11 74.56 30.80 66.06 55.09 54.34 67.19 24.09 53.07 34.81 51.11
[…]
@@ 6.5 Decoding Speed on Pruned MoE Layer
6.6 Integrity of Micro-Experts
A central insight of our method is that jointly compressing multiple matrices preserves the functional integrity of micro-experts. Figure 5 compares three different methods. All share the same FFN structure: the input $\mathbf{x}$ is transformed by $\mathbf{W}^{\mathrm{up}}$ and $\mathbf{W}^{\mathrm{gate}}$ into intermediate activations, followed by $\mathbf{W}^{\mathrm{down}}$ to produce $\mathbf{y}$. As highlighted in the red dashed box, each row of $\mathbf{W}^{\mathrm{up}}$ maps uniquely to a column of $\mathbf{W}^{\mathrm{down}}$, forming a one-to-one micro-expert. For visualization, $\mathbf{W}^{\mathrm{up}}$ is shown transposed and micro-experts are energy-ranked from left to right.
Camera-P prunes low-energy micro-experts while keeping the remains intact. Camera-Q assigns higher precision to higher-energy micro-experts, enforcing uniform precision within each. In contrast, Camera-Q slices $\mathbf{W}^{\mathrm{up}}$ and $\mathbf{W}^{\mathrm{gate}}$ along an orthogonal dimension, quantizing each weight with multiple precisions. Like many existing mixed-precision approaches, Camera-Q operates on individual matrices and allocates bits along the input dimension (e.g., using $\mathbf{H}=\mathbf{XX}^{\mathrm{T}}$), overlooking cross-matrix expert structure. This often gives high-energy micro-experts inconsistent or insufficient precision. Our experiments show that enforcing precision consistency within each micro-expert, as in Camera-Q, yields substantially better performance.
6.7 Application on Dense Model
Camera investigates redundancy that spans across matrices within MoE layers. After applying cross-matrix compression, the resulting MoE layers can still be further compressed using single-matrix methods such as Wanda (wanda2024). Moreover, our approach can also be directly applied to structured pruning or mixed-precision quantization in dense-architecture models. As an illustrative example, we apply Camera-P to prune 20% of the FFN layers in the openPangu-7B model (pangu2025) running on Ascend 910B. On top of this, we further apply Wanda to perform an additional 20% unstructured pruning. The resulting performance is summarized in Table 5. These results demonstrate that expert-level pruning and intra-matrix pruning can be seamlessly combined to achieve higher overall pruning ratios while maintaining lossless performance.
@@ 7 Conclusion
Acknowledgement
We gratefully acknowledge the support of the National Natural Science Foundation of China (NSFC) via grants 62236004, 62206078 and 62476073. Additional gratitude is extended to our collaborators and colleagues for their insightful discussions during the development of this work.
Appendix
According to the AAAI Proceedings policy, appendices and supplementary materials are not included in the published proceedings. We therefore provide the complete version of the paper, including all appendices, on arXiv for reference of readers. Please refer to https://arxiv.org/abs/2508.02322 for the complete appendix.
@@ A.1 Derivation of Micro-Expert
$\mathbf{y}$ $=\sum_{i}^{N_{E}}A_{i}\cdot\sum_{j}^{d_{\mathrm{ff}}}\sigma\left(\mathbf{W}_{i}^{\mathrm{gate}}[j,:]\mathbf{x}\right)\cdot\mathbf{W}_{i}^{\mathrm{up}}[j,:]\mathbf{x}\cdot\mathbf{W}_{i}^{\mathrm{down}}[:,j].$
@@
$\mathbf{y}$ $=\sum_{i}^{N_{E}\times d_{\mathrm{ff}}}A_{i}\cdot\sigma\left(\mathbf{w}_{i}^{\mathrm{gate}}\mathbf{x}\right)\cdot\mathbf{w}_{i}^{\mathrm{up}}\mathbf{x}\cdot\mathbf{w}^{\mathrm{down}}.$
We separate the scalar terms in the above equation and denote them as $\phi_{i}=A_{i}\cdot\sigma\left(\mathbf{w}_{i}^{\mathrm{gate}}\mathbf{x}\right)\cdot\mathbf{w}_{i}^{\mathrm{up}}\mathbf{x}$. Hence, we have:
@@ A.3 Supplementary Micro-Expert Distributions
Figure 7: Distributions of micro-experts within each expert based on global ranking from Camera on Deepseek-MoE-16B. We list all 66 experts, where ‘S0/S1’ denotes the shared experts, and the rest are non-shared common experts. For 7(c), we present the results of the pruned model at $\lambda=40\%$, while the other subfigures correspond to $\lambda=20\%$.
@@
In Section 6.1 of the main text, we presented the micro-expert analysis of layer 12 in the Deepseek-MoE-16B model with a overall pruning ratio of $\lambda=20\%$. Here, we further supplement the analysis with results from layers 3 and 28, as shown in Figures 7(a) and 7(b). First, we observe that the importance of the first two shared experts, S0 and S1, is quite similar across layers, though not exactly the same. In contrast, the importance of the remaining experts varies significantly. Additionally, we also find that the shared experts tend to be less important in the lower layers (e.g., layers 2 and 3), as illustrated in Figure 7(a). Finally, we highlight a seemingly important property of the Camera algorithm: the micro-expert analysis results—namely, the importance rankings—remain largely consistent across different values of $\lambda$. As shown in Figure 7(c), when $\lambda=40\%$, the analysis results closely resemble those in Figure 2.
@@
As shown in Figure 8, the micro-expert distribution of the Qwen3-30B-A3B model is similar to that of Deepseek-MoE-16B, exhibiting large disparities in expert importance. Unlike Deepseek-MoE-16B, Qwen3-30B-A3B does not employ shared experts—all 128 experts are treated equally. Moreover, we observe that Qwen3-30B-A3B appears to have a higher proportion of unimportant experts. As shown in Figures 8(b) and 8(c), more expert box plots are concentrated near the top. We speculate that this may result from insufficient weight utilization, suggesting potential for continued training and further knowledge injection.